A survey group of 6 students are chosen at random from 25 elementary students and 13 high school students. Determine the number of ways that at least 2 of the students chosen are elementary students.
Answer by Boreal(15235)

Find a way where only 1 elementary student is chosen
That is 25C1*13C5=25*1287=32175
Find a way where 0 is chosen
That is 13C6=1716
Now find the way 6 can be chosen from 38.
That is 2760681 total choices
subtract the above 2 sum or 33891
That leaves 2,726,790 ways at least 2.